4t-t=6 + t <---- how do u do this?!?!?!
Answer by checkley75(3666)

4t-t=6+t you need to combine the t's on the left side of the equation
4t-t-t=6 now combine the t's
2t=6 now divide both sides by 2
2t/2=6/2 now cancel/reduce these fractions
t=3 answer.
proof
4*3-3=6+3
12-3=9
9=9